Identify the limiting reactant when 1.7 g of sodium reacts with 2.6 l of chlorine gas at stp to produce sodium chloride?

1
Expert's answer
2021-11-22T17:19:04-0500

Solution:

Calculate the moles of each reactant:

The molar mass of sodium (Na) is 23 g/mol

Hence,

(1.7 g Na) × (1 mol Na / 23 g Na) = 0.074 mol Na


At STP, one mole of any gas occupies a volume of 22.4 L.

Hence,

(2.6 L Cl2) × (1 mol Cl2 / 22.4 L Cl2) = 0.116 mol Cl2


Balanced chemical equation:

2Na(s) + Cl2(g) → 2NaCl(s)

According to stoichiometry:

2 mol of Na reacts with 1 mol of Cl2

Thus, 0.074 mol of Na reacts with:

(0.074 mol Na) × (1 mol Cl2 / 2 mol Na) = 0.037 mol Cl2

However, initially there is 0.116 mol of Cl2 (according to the task).

Therefore, Na acts as limiting reactant and Cl2 is excess reactant.


Answer: Sodium (Na) acts as limiting reactant.